Frank wants a cleaning solution that is 40% acid. He has a large supply of both 25% and 50% solution. How many liters of each does he need to mix to make 30liters of the desired solution?
Answer by mananth(16949)

percent ---------------- quantity
25 ---------------- x
50 ---------------- 30-x
40.00% ---------------- 30
..............
25x+50(30-x)=40 *30
25x+1500-50x=1200
25x-50x =1200-1500
-25x=-300
/-25
x=12liters of 25.00%
18 liters of 50.00%